# Manifolds with kinks and the asymptotic behavior of the graph Laplacian operator with Gaussian kernel

### v2

@@ 4.3. Euclidean calculation
Recall that $c_{\ell}:=\Gamma((\ell+1)/2)/2$ for any positive integer $\ell$, where $\Gamma$ is the usual Gamma function. We prove the following result where we use the notation introduced in (13).
@@
Let $p\in\mathcal{C}^{2}_{\geq 0}(\mathbb{B}^{d})$ and $f\in\mathcal{C}^{3}(\mathbb{B}^{d})$ be such that
(20) $C_{p,f}:=\left(\max_{0\leq j\leq 2}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(j)}_{\xi}p\|_{op}+\max_{1\leq i\leq 3}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(i)}_{\xi}f\|_{op}\right)<+\infty.$
@@
$L_{t}^{\mathcal{C},\eta}f(0_{d})=\frac{1}{t}\int_{\mathbb{B}^{d}_{t^{\eta}}\cap\mathcal{C}}e^{-\frac{\|y\|^{2}}{t}}(f(0_{d})-f(y))p(y)\mathop{}\!\mathrm{d}y.$
@@
(21) $L_{t}^{\mathcal{C},\eta}f(0_{d})$ $=-\frac{c_{d}}{\sqrt{t}}p(0_{d})\partial_{v_{\mathcal{C}}}f(0_{d})-c_{d+1}\bigg{(}p(0_{d})A_{\mathcal{C}}f(0_{d})+[p,f]_{\mathcal{C}}(0_{d})\bigg{)}+O(\sqrt{t}),$
@@
$\partial_{v_{\mathcal{C}}}f(0_{d}):=d_{0_{d}}f(v_{\mathcal{C}})\quad\text{with}\quad v_{\mathcal{C}}:=\int_{S^{g}\mathcal{C}}\theta\mathop{}\!\mathrm{d}\sigma(\theta),$
$A_{\mathcal{C}}f(0_{d}):=\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}f\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta)\qquad\text{and}\qquad[p,f]_{\mathcal{C}}(x):=\frac{1}{2}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}f(\theta)\mathop{}\!\mathrm{d}_{0_{d}}p(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
By asymptotic property of the incomplete Gamma function (see e.g. [Pin20])
@@
The result follows from the fact that $C_{g}t^{a(m+d-2)}e^{-t^{2a}}=O(\sqrt{t})$ as $t\downarrow 0$. ∎
@@
$L_{t}^{\mathcal{C},\eta}f(0_{d})$ $=-t^{-\frac{1}{2}}p(0_{d})\mathop{}\!\mathrm{d}_{0_{d}}f\left(\int_{\mathcal{C}}e^{-\|z\|^{2}}z\mathop{}\!\mathrm{d}z\right)-\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(z)\mathop{}\!\mathrm{d}_{0_{d}}p(z)\mathop{}\!\mathrm{d}z$
(22) $\phantom{=}-\frac{p(0_{d})}{2}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}f\left(\int_{\mathcal{C}}e^{-\|z\|^{2}}z^{(2)}\mathop{}\!\mathrm{d}z\right)+O(\sqrt{t}).$
@@
$f(0_{d})-f(y)=-\mathop{}\!\mathrm{d}_{0_{d}}f(y)-\frac{1}{2}\mathop{}\!\mathrm{d}^{(2)}_{0_{d}}f(y^{(2)})-\frac{1}{6}\mathop{}\!\mathrm{d}^{(3)}_{\xi_{y}}f(y^{(3)}),$
$p(y)=p(0_{d})+\mathop{}\!\mathrm{d}_{0_{d}}p(y)+\frac{1}{2}\mathop{}\!\mathrm{d}^{(2)}_{\zeta_{y}}p(y^{(2)}).$
@@
$|I_{t}(i,j)|$ $\leq\frac{1}{t^{\frac{d}{2}+1}}\int_{\mathcal{C}\cap\mathbb{B}^{d}_{t^{\eta}}}e^{-\frac{|y|^{2}}{t}}|h_{i,j}(y)|\mathop{}\!\mathrm{d}y\leq\frac{C_{p,f}}{t^{\frac{d}{2}+1}}\int_{\mathcal{C}\cap\mathbb{B}^{d}_{t^{\eta}}}e^{-\frac{|y|^{2}}{t}}|y|^{i+j}\mathop{}\!\mathrm{d}y$
$=C_{p,f}t^{\frac{i+j}{2}-1}\int_{\mathcal{C}\cap\mathbb{B}^{d}_{t^{\eta-1/2}}}e^{-\|z\|^{2}}\|z\|^{i+j}\mathop{}\!\mathrm{d}z\leq C_{p,f}C_{\mathcal{C},i+j}t^{\frac{i+j}{2}-1}$
@@
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{1,0}(z)\mathop{}\!\mathrm{d}z$ $=p(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{1,1}(z)\mathop{}\!\mathrm{d}z$ $=\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(z)\mathop{}\!\mathrm{d}_{0_{d}}p(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{2,0}(z)\mathop{}\!\mathrm{d}z$ $=p(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}f(z^{(2)})\mathop{}\!\mathrm{d}z.$
@@
$\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(z)\mathop{}\!\mathrm{d}_{0_{d}}p(z)\mathop{}\!\mathrm{d}z$ $=\int_{0}^{+\infty}\int_{S^{g}\mathcal{C}}e^{-\rho^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(\rho\theta)\mathop{}\!\mathrm{d}_{0_{d}}p(\rho\theta)\rho^{d-1}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\sigma(\theta)$
$=\int_{0}^{+\infty}\int_{S^{g}\mathcal{C}}e^{-\rho^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(\theta)\mathop{}\!\mathrm{d}_{0_{d}}p(\theta)\rho^{d+1}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\sigma(\theta)$
$=\left(\int_{0}^{+\infty}e^{-\rho^{2}}\rho^{d+1}\mathop{}\!\mathrm{d}\rho\right)\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}f(\theta)\mathop{}\!\mathrm{d}_{0_{d}}p(\theta)\mathop{}\!\mathrm{d}\sigma(\theta)$
$=c_{d+1}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}f(\theta)\mathop{}\!\mathrm{d}_{0_{d}}p(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
The previous proof may be easily modified to get the following: for any integer $N\geq 2$, let $p\in\mathcal{C}^{N}_{\geq 0}(\mathbb{B}^{d})$ and $f\in\mathcal{C}^{N+1}(\mathbb{B}^{d})$ be such that
$\max_{0\leq j\leq N}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(j)}_{\xi}p\|_{op}+\max_{1\leq i\leq N+1}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(i)}_{\xi}f\|_{op}<+\infty.$
@@
$L_{t}^{\mathcal{C},\eta}f(0_{d})=\sum_{i=1}^{N+1}\sum_{j=0}^{N}\frac{t^{\frac{i+j}{2}-1}c_{d+i+j-1}}{i!j!}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}^{(i)}f(\theta^{(i)})\mathop{}\!\mathrm{d}_{0_{d}}^{(j)}p(\theta^{(j)})\mathop{}\!\mathrm{d}\sigma(\theta)+O(t^{\frac{N-1}{2}}).$
Moreover, it should be pointed out that if $p(0_{d})=0$, then $L_{t}^{\mathcal{C},\eta}f(0_{d})$ always converges to
$\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}f(z)\mathop{}\!\mathrm{d}_{0_{d}}p(z)\mathop{}\!\mathrm{d}z.$
@@ 4.4. Conclusion
As for the other two cases, it follows from (5) for interior points, and from Proposition 2.11 and Proposition 2.13 for LCDD border points, that we can replace $\tilde{F}_{0_{d}}(\bar{\Omega})$ by $I_{0_{d}}\bar{\Omega}$ in the previous domain of integration. Then we apply Proposition 4.4 to obtain that
@@ 5. Asymptotic behavior of the extrinsic Gaussian Operator
In this section, we prove Theorem 1.2. Let $M\subset\mathbb{R}^{D}$ be a smooth $d$-dimensional submanifold with kinks endowed with the Riemannian metric $g$ induced by the Euclidean scalar product of $\mathbb{R}^{D}$. Recall that the extrinsic Gaussian operator at time $t>0$ associated with a $\mathcal{C}^{2}$ density $p$ on $M$ is given by
@@
Following [CL06], for a generic $y\in\mathbb{B}_{\epsilon}^{D}\cap M$ we set
@@
It follows line by line from [CL06, Appendix B, Lemma 7], with the assumption that $\left\lVert y-x\right\rVert<t^{\eta}$ instead of $t^{1/2}$ like they did, which and whose implications are both indeed weaker, since $t^{\eta}>t^{1/2}$ for small $t>0$. ∎

### v3

@@ 4.3. Euclidean calculation
With a view to apply it to the expression of $I(t)$ established in (35), we prove the next Euclidean result. Recall that $c_{\ell}:=\Gamma((\ell+1)/2)/2$ for any positive integer $\ell$, where $\Gamma$ is the Gamma function.
@@
Let $q\in\mathcal{C}^{2}_{\geq 0}(\mathbb{B}^{d})$ and $h\in\mathcal{C}^{3}(\mathbb{B}^{d})$ be such that
(27) $C_{q,h}:=\left(\max_{0\leq j\leq 2}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(j)}_{\xi}q\|_{op}+\max_{1\leq i\leq 3}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(i)}_{\xi}h\|_{op}\right)<+\infty.$
@@
$L_{t}^{\mathcal{C},\eta}h(0_{d})=\frac{1}{t}\int_{\mathbb{B}^{d}_{t^{\eta}}\cap\mathcal{C}}e^{-\frac{\|y\|^{2}}{t}}(h(0_{d})-h(y))q(y)\mathop{}\!\mathrm{d}y.$
@@
(28) $L_{t}^{\mathcal{C},\eta}h(0_{d})$ $=-\frac{c_{d}}{\sqrt{t}}q(0_{d})\partial_{v_{\mathcal{C}}}h(0_{d})-c_{d+1}\bigg(q(0_{d})A_{\mathcal{C}}h(0_{d})+[q,h]_{\mathcal{C}}(0_{d})\bigg)+O(\sqrt{t}),$
@@
$\partial_{v_{\mathcal{C}}}h(0_{d}):=d_{0_{d}}h(v_{\mathcal{C}})\quad\text{with}\quad v_{\mathcal{C}}:=\int_{S^{g}\mathcal{C}}\theta\mathop{}\!\mathrm{d}\sigma(\theta),$
$A_{\mathcal{C}}h(0_{d}):=\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h\left(\theta^{(2)}\right)\mathop{}\!\mathrm{d}\sigma(\theta)\qquad\text{and}\qquad[q,h]_{\mathcal{C}}(x):=\frac{1}{2}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}h(\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
By asymptotic property of the incomplete Gamma function (see e.g. [Pinelis])
@@
The result follows from the fact that $C_{h}t^{a(m+d-2)}e^{-t^{2a}}=O(\sqrt{t})$ as $t\downarrow 0$. ∎
@@
$L_{t}^{\mathcal{C},\eta}h(0_{d})$ $=-t^{-\frac{1}{2}}q(0_{d})\mathop{}\!\mathrm{d}_{0_{d}}h\left(\int_{\mathcal{C}}e^{-\|z\|^{2}}z\mathop{}\!\mathrm{d}z\right)-\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}_{0_{d}}q(z)\mathop{}\!\mathrm{d}z$
(29) $\phantom{=}-\frac{q(0_{d})}{2}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h\left(\int_{\mathcal{C}}e^{-\|z\|^{2}}z^{(2)}\mathop{}\!\mathrm{d}z\right)+O(\sqrt{t}).$
@@
$h(0_{d})-h(y)=-\mathop{}\!\mathrm{d}_{0_{d}}h(y)-\frac{1}{2}\mathop{}\!\mathrm{d}^{(2)}_{0_{d}}h(y^{(2)})-\frac{1}{6}\mathop{}\!\mathrm{d}^{(3)}_{\xi_{y}}h(y^{(3)}),$
$q(y)=q(0_{d})+\mathop{}\!\mathrm{d}_{0_{d}}q(y)+\frac{1}{2}\mathop{}\!\mathrm{d}^{(2)}_{\zeta_{y}}q(y^{(2)}).$
@@
$|I_{t}(i,j)|$ $\leq\frac{1}{t^{\frac{d}{2}+1}}\int_{\mathcal{C}\cap\mathbb{B}^{d}_{t^{\eta}}}e^{-\frac{|y|^{2}}{t}}|h_{i,j}(y)|\mathop{}\!\mathrm{d}y\leq\frac{C_{q,h}}{t^{\frac{d}{2}+1}}\int_{\mathcal{C}\cap\mathbb{B}^{d}_{t^{\eta}}}e^{-\frac{|y|^{2}}{t}}|y|^{i+j}\mathop{}\!\mathrm{d}y$
$=C_{q,h}t^{\frac{i+j}{2}-1}\int_{\mathcal{C}\cap\mathbb{B}^{d}_{t^{\eta-1/2}}}e^{-\|z\|^{2}}\|z\|^{i+j}\mathop{}\!\mathrm{d}z\leq C_{q,h}C_{\mathcal{C},i+j}t^{\frac{i+j}{2}-1}$
@@
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{1,0}(z)\mathop{}\!\mathrm{d}z$ $=q(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{1,1}(z)\mathop{}\!\mathrm{d}z$ $=\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}_{0_{d}}q(z)\mathop{}\!\mathrm{d}z$
$\int_{\mathcal{C}}e^{-\|z\|^{2}}g_{2,0}(z)\mathop{}\!\mathrm{d}z$ $=q(0_{d})\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}^{(2)}h(z^{(2)})\mathop{}\!\mathrm{d}z.$
@@
$\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}_{0_{d}}q(z)\mathop{}\!\mathrm{d}z$ $=\int_{0}^{+\infty}\int_{S^{g}\mathcal{C}}e^{-\rho^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(\rho\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\rho\theta)\rho^{d-1}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\sigma(\theta)$
$=\int_{0}^{+\infty}\int_{S^{g}\mathcal{C}}e^{-\rho^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\theta)\rho^{d+1}\mathop{}\!\mathrm{d}\rho\mathop{}\!\mathrm{d}\sigma(\theta)$
$=\left(\int_{0}^{+\infty}e^{-\rho^{2}}\rho^{d+1}\mathop{}\!\mathrm{d}\rho\right)\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}h(\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\theta)\mathop{}\!\mathrm{d}\sigma(\theta)$
$=c_{d+1}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}h(\theta)\mathop{}\!\mathrm{d}_{0_{d}}q(\theta)\mathop{}\!\mathrm{d}\sigma(\theta).$
@@
The previous proof may be easily modified to get the following: for any integer $N\geq 2$, let $q\in\mathcal{C}^{N}_{\geq 0}(\mathbb{B}^{d})$ and $h\in\mathcal{C}^{N+1}(\mathbb{B}^{d})$ be such that
$\max_{0\leq j\leq N}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(j)}_{\xi}q\|_{op}+\max_{1\leq i\leq N+1}\sup_{\xi\in\mathbb{B}^{d}}\|\mathop{}\!\mathrm{d}^{(i)}_{\xi}h\|_{op}<+\infty.$
@@
$L_{t}^{\mathcal{C},\eta}h(0_{d})=\sum_{i=1}^{N+1}\sum_{j=0}^{N}\frac{t^{\frac{i+j}{2}-1}c_{d+i+j-1}}{i!j!}\int_{S^{g}\mathcal{C}}\mathop{}\!\mathrm{d}_{0_{d}}^{(i)}h(\theta^{(i)})\mathop{}\!\mathrm{d}_{0_{d}}^{(j)}q(\theta^{(j)})\mathop{}\!\mathrm{d}\sigma(\theta)+O(t^{\frac{N-1}{2}}).$
Moreover, it should be pointed out that if $q(0_{d})=0$, then $L_{t}^{\mathcal{C},\eta}h(0_{d})$ always converges to
$\int_{\mathcal{C}}e^{-\|z\|^{2}}\mathop{}\!\mathrm{d}_{0_{d}}h(z)\mathop{}\!\mathrm{d}_{0_{d}}q(z)\mathop{}\!\mathrm{d}z.$
@@ 4.4. Conclusion
As for the other two cases, it follows from (6) for interior points, and from Proposition 2.11 and Proposition 2.13 for LCDD border points, that we can replace $\tilde{F}_{0_{d}}(\bar{\Omega})$ by $I_{0_{d}}\bar{\Omega}$ in the previous domain of integration. Then we apply Proposition 4.3 with $\mathcal{C}=I_{0_{d}}\bar{\Omega}$ to obtain that
@@ 5. Asymptotic behavior of the extrinsic Gaussian Operator
In this section, we prove Theorem 1.2 without the refined estimates on the error term $\mathrm{Err}(t)$. See Section 6 for these refined estimates. Let $M\subset\mathbb{R}^{D}$ be a smooth $d$-dimensional submanifold with kinks endowed with the Riemannian metric $g$ induced by the Euclidean scalar product of $\mathbb{R}^{D}$. Recall that the extrinsic Gaussian operator at time $t>0$ associated with a $\mathcal{C}^{2}$ density $p$ on $M$ is given by
@@
Following [CoifmanLafon2006], for a generic $y\in\mathbb{B}_{\epsilon}^{D}\cap M$ we set
@@
It follows line by line from [CoifmanLafon2006, Appendix B, Lemma 7], with the assumption that $\left\lVert y-x\right\rVert<t^{\eta}$ instead of $t^{1/2}$ like they did, which and whose implications are both indeed weaker, since $t^{\eta}>t^{1/2}$ for small $t>0$. ∎